# A Gaussian integral formula for the Hermite polynomials: Combinatorics, Asymptotics and Applications

### v1

@@ 3.4 Three approximations to $H\!\!\operatorname{e}_{n}(x)$ when $x=O(\sqrt{n})$
It is also natural to obtain asymptotics for $H\!\!\operatorname{e}_{n}(x)$ when $x$ grows proportionally to $\sqrt{n}$, i.e. when $x=2a\sqrt{n}$ for some constant $a$. The results here are equivalent to the so-called Plancherel–Rotach asymptotics for the Hermite polynomials (see e.g. Theorem 8.22.9 of [Sze75]). However, our results are stated and proven without any reference to $n!$ (which arises from the Cauchy integral formula in the traditional Plancherel–Rotach setup).
@@
  • 2. Large-$x$ regime $\mathbf{x>2\sqrt{n}}$: For fixed $a>1$, write $a=\cosh(\gamma)$ with $\gamma\geq 0$ and set $x=2a\sqrt{n}$. We obtain:
@@
$\widetilde{H}_{n}\left(a,\frac{1}{4n}\right)=\mathbb{E}_{W\sim\mathcal{N}(0,\frac{1}{n})}\left[\left(a+i\frac{W}{2}\right)^{n}\right]=\sqrt{\frac{n}{2\pi}}\intop_{-\infty}^{\infty}(a+i\frac{w}{2})^{n}e^{-\frac{1}{2}nw^{2}}\mathrm{d}w$
@@
This leads to constant phase curves visualized in Figure 5 . Then, it remains to apply the Laplace method to the integral to get the limiting behaviour of $\widetilde{H}(a,\frac{1}{4n})$ in each of the cases. The details of the this are explained in Lemmas 3.9, 3.8 and the analogous requited limit for $\widetilde{H}_{n}(1+\frac{1}{2}un^{-2/4},\frac{1}{4n})$ in Lemma 3.12
@@ 3.4.2 Proof of edge asymptotics
Let $\langle$ be the union of the line segments $\left\{te^{i\pi/3}\colon t\geq 0\right\}$ and $\left\{te^{-i\pi/3}\colon t\geq 0\right\}$, that is the union of the two infinite rays in the complex plane emanating from the origin that make angles $\pi/3$ and $\pi/3$ with the real axis (see Figure 6). For $u\in\mathbb{C}$ the Airy function is defined by the complex contour integral
@@
Hence, this part of the integral is going to zero exponentially fast and $\widetilde{H}_{n}$ is well approximated by an integral over only the part near the critical point $\Gamma^{\ast}_{\leq}$. Notice also that since $|z|\leq n^{\epsilon}$ on $\Gamma^{\ast}_{\leq}$, so we can control the error term in our series expansion 57 by $O((w-i)^{4})=O(\frac{z^{4}}{n^{4/3}})=O(n^{4\epsilon-4/3})$. Hence, as long as $4\epsilon-1/3<0$, even when multiplying by $n$ we have a good approximation in this region:
@@
4 Applications to the oscillator wave functions $\Psi_{n}$ and Hermite kernel $K_{n}$
@@ 4.1 Definitions of $\Psi_{n}$ and $K_{n}$
The asymptotics of Section 3.4 reveals that there is a critical phase transition in the Hermite polynomials $H\!\!\operatorname{e}_{n}(x)$ at the location $x=\pm 2\sqrt{n}$ . For large $x$ values, $|x|>2\sqrt{n}$, the growth of the Hermite polynomials is overwhelmed by the Gaussian density, and we see that the $\Psi_{n}(x)$ decays exponentially fast to zero in this region. We make this precise as follows
@@
Define the function $f(a)=-\text{arccosh}(a)-\exp(-2\text{arccosh}(a))/2+a^{2}-\frac{1}{2}$. Then $f(a)>0$ for all $a>1$, $f$ is increasing in $a$, and for any fixed $a$ with $a>1$, we have:
$\lim_{n\to\infty}\frac{|\Psi_{n}(2\sqrt{n}a)|}{\exp\left(-nf(a)\right)}=0.$
In other words, $\Psi_{n}(2\sqrt{n}a)\ll\exp(-nf(a))$ goes to zero exponentially fast as $n\to\infty$.
@@
By (43), setting $\gamma=\text{arccosh}(a)$, $a=\cosh(\gamma)$ with $\gamma>0$, we have the limit:
$|\Psi_{n}(2\sqrt{n}a)|$ $=\frac{1}{\sqrt{n!}}|H\!\!\operatorname{e}_{n}(2\sqrt{n}a)|\sqrt{\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}(2\sqrt{n}a)^{2}}}$
$=\frac{1}{{2\pi}^{1/4}\sqrt{n!}}n^{n/2}e^{n(\gamma+e^{-2\gamma}/2)}\,\frac{e^{\gamma/2}}{\sqrt{2\sinh(\gamma)}}e^{-n\cosh(\gamma)^{2}}\left(1+o(1)\right)$
By Stirling’s approximation, Example 3.4 , $n!=\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}(1+o(1))$, so we have after some cancellation
$|\Psi_{n}(2\sqrt{n}a)|$ $=\frac{1}{\sqrt{2\pi}}\frac{1}{n^{1/4}}e^{n(\gamma+e^{-2\gamma}/2-\cosh(\gamma)^{2}+\frac{1}{2})}\,\frac{e^{\gamma/2}}{\sqrt{2\sinh(\gamma)}}\left(1+o(1)\right)$
$=\frac{1}{\sqrt{2\pi}}\frac{1}{n^{1/4}}e^{-nf(a)}\,\frac{e^{\gamma/2}}{\sqrt{2\sinh(\gamma)}}\left(1+o(1)\right)$
Elementary calculus shows that $f(a)>0$ for $a>1$ (in fact, one can further show that $f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0$ and all higher derivatives are positive, $f^{(k)}(0)>0$). The result now follows since $n^{1/4}\to\infty$ as $n\to\infty$. ∎
@@ 4.3 Cramèr’s bound for the global maximum
The function $\Psi_{n}(x)$ goes to zero exponentially fast outside of $x=[-2\sqrt{n},2\sqrt{n}]$, and the asymptotics of Section 3.4 suggest it should achieve its global maximum $\max_{x\in\mathbb{R}}|\Psi_{n}(x)|$ somewhere near the edge of this region $x=\pm 2\sqrt{n}$. However, since all of the asymptotics are proven pointwise, a different approach is needed in order to draw precise conclusions about the actual maximum. In [Ind61], a clever elementary analysis of the zeros of $\Psi^{\prime}_{n}$ reveals indeed that the maximum occurs at the closest zero of $\Psi_{n}^{\prime}$ to the edge $2\sqrt{n}$ and finds a bound for the maximum $\max_{x\in\mathbb{R}}|\Psi_{n}(x)|$. This bound is sometimes called Cramèr’s bound. We state this result here for later use.
@@
The proof is given in [Ind61]. Note that the function called $E_{n}(x)$ there is related to $\Psi_{n}(x)$ by $\Psi_{n}(x)=\frac{1}{2^{1/4}}E_{n}(x/\sqrt{2})$ . ∎
@@ 4.4 Arcsine law limit for $\Psi_{n}^{2}$ and the limit of the quantum harmonic oscillator
By Proposition 4.3 $\Psi_{n}(\sqrt{n}x)\to 0$ for $|x|>2$ exponentially fast. To see the convergence to the function $\frac{1}{\pi\sqrt{4-x^{2}}}$ for $x\in[-2,2]$, we use the approximation (42) with $a=x/2$ to get
@@
The statement of the Schroedinger equation is exactly as in [Gri18], where the quantization for the energy is derived and it is shown that the solution is $\psi_{n}(x)=\left(\frac{m\omega}{\pi\hbar}\right)^{1/4}\frac{1}{\sqrt{2^{n}n!}}H_{n}(\xi)e^{-\xi^{2}/2}$ where $H_{n}$ are the physicists Hermite polynomials ($H_{1}(x)=2x,H_{2}(x)=4x^{2}-2$, etc) and $\xi=\sqrt{m\omega/\hbar}x$ is rescaled space. This is identical to our stated version by using the translation between physicists and probabilist $H_{n}(x)=2^{n/2}H\!\!\operatorname{e}_{n}(\sqrt{2}x)$ and noting that $cx=\sqrt{2}\xi$ by its definition. (One can also directly verify using the change of variable Section 6 that this solves the stated Schroedinger PDE)
@@ 4.5 Semi-circle law limit for $K_{n}$
The first statement follows by the simple change of variable $y=cx$ to establish that for any distribution $\rho$ and any constant $c$, one has:
@@
By the convergence of $\sqrt{n}\Psi_{n}(\sqrt{n}x)^{2}=\frac{\sqrt{n}}{\sqrt{2\pi}n!}H\!\!\operatorname{e}_{n}(\sqrt{n}x)^{2}e^{-\frac{1}{2}nx^{2}}$ to the arcsine law of Proposition 4.5, the compact support of the arcinse distribution on [2,2][-2,2] and the exponential decay of Propositon 4.3, we conclude that the moments of $\sqrt{n}\Psi_{n}(\sqrt{n}x)^{2}$ converge to the moments of the arcsine law.
Now, because the semi-circle law is the corresponding integral of the arcsine law from 63, by application of Lemma 4.9 we consequently see that the moments of $\sqrt{n}^{-1}K_{n}(\sqrt{n}x,\sqrt{n}x)$ converge to the moments of the semi-circle law. Since the semi-circle law is compactly supported on [2,2][-2,2], it is the unique distribution with these moments, and therefore the moment convergence implies the convergence in distribution (See e.g. Section 3.3.5 of [Dur10] on the moment problem and in particular Theorem 3.3.12) ∎
@@
The connection between the semi-circle law and the arcsine law has been noted in many other contexts too. In [HV19] the convergence is proven using random graphs. In [Led04], the convergence to the arcsine law is shown using differential equations. It is subsequently noted there that the mixture converges to $\sqrt{U}\xi_{arcsine}$ where $\xi_{arcsine}$ has the arcsine law, and $U$ is an independent uniform random variable. One can verify that $\sqrt{U}\xi_{arcsine}$ is distributed according to the semicircle law.

### v2

@@ 3.4 Three approximations to $H\!\!\operatorname{e}_{n}(x)$ when $x=O(\sqrt{n})$
It is also natural to obtain asymptotics for $H\!\!\operatorname{e}_{n}(x)$ when $x$ grows proportionally to $\sqrt{n}$, i.e. when $x=2a\sqrt{n}$ for some constant $a$. The results here are equivalent to the so-called Plancherel–Rotach asymptotics for the Hermite polynomials (see e.g. Theorem 8.22.9 of [Szego1975]). However, our results are stated and proven without any reference to $n!$ (which arises from the Cauchy integral formula in the traditional Plancherel–Rotach setup).
@@
  • 2. Large-$x$ regime $\bm{x>2\sqrt{n}}$: For fixed $a>1$, write $a=\cosh(\gamma)$ with $\gamma\geq 0$ and set $x=2a\sqrt{n}$. We obtain:
@@
$\widetilde{H}_{n}\left(a,\frac{1}{4n}\right)=\mathbb{E}_{W\sim\mathcal{N}(0,\frac{1}{n})}\left[\left(a+i\frac{W}{2}\right)^{n}\right]=\sqrt{\frac{n}{2\pi}}\intop_{-\infty}^{\infty}\left(a+i\frac{w}{2}\right)^{n}e^{-\frac{1}{2}nw^{2}}\mathrm{d}w$
@@
This leads to constant phase curves visualized in Figure 5 . Then, it remains to apply the Laplace method to the integral to get the limiting behaviour of $\widetilde{H}(a,\frac{1}{4n})$ in each of the cases. The details of the this are explained in Lemmas 3.10, 3.9 and the analogous requited limit for $\widetilde{H}_{n}(1+\frac{1}{2}un^{-2/4},\frac{1}{4n})$ in Lemma 3.13
@@ 3.4.2 Proof of edge asymptotics
Let $\langle$ denote the union of the line segments $\left\{te^{i\pi/3}\colon t\geq 0\right\}$ and $\left\{te^{-i\pi/3}\colon t\geq 0\right\}$, that is the union of the two infinite rays in the complex plane emanating from the origin that make angles $\pi/3$ and $\pi/3$ with the real axis (see Figure 6). For $u\in\mathbb{C}$, the Airy function is defined by the complex contour integral
@@
Hence, this part of the integral is going to zero exponentially fast and $\widetilde{H}_{n}$ is well approximated by an integral over only the part near the critical point $\Gamma^{\ast}_{\leq}$. Notice also that since $|z|\leq n^{\epsilon}$ on $\Gamma^{\ast}_{\leq}$, so we can control the error term in our series expansion (60) by $O((w-i)^{4})=O(\frac{z^{4}}{n^{4/3}})=O(n^{4\epsilon-4/3})$. Hence, as long as $4\epsilon-1/3<0$, even when multiplying by $n$ we have a good approximation in this region:
@@
3.5 Exponentially fast decay for $a>1$
The asymptotics of Section 3.4 reveal a critical phase transition in the Hermite polynomials $H\!\!\operatorname{e}_{n}(x)$ at the location $x=\pm 2\sqrt{n}$ . For large $x$ values, $|x|>2\sqrt{n}$, the growth of the Hermite polynomials is overwhelmed by the Gaussian density. Here we obtain an upper bound that holds for all $n$.
Proposition 3.14.
For all $a>1$ and all $n\in\mathbb{N}$,
$\left|H\!\!\operatorname{e}_{n}(2\sqrt{n}a)\right|$ $\leq\frac{n^{n/2}}{\sqrt{2(1-a^{2}+a\sqrt{a^{2}-1}})}\,e^{n(a^{2}-a\sqrt{a^{2}-1}-\frac{1}{2}+\text{arccosh}(a))}$
Remark 3.15.
Note that the exponential growth rate in Proposition prop:bound-uniform-H matches that of the largex-x regime of Theorem 3.7: if $a=\cosh(\gamma)$ then
$e^{-2\gamma}$ $=\frac{1}{(a+\sqrt{a^{2}-1})^{2}}=(a-\sqrt{a^{2}-1})^{2}=2a^{2}-2a\sqrt{a^{2}-1}-1$
and so
$e^{n(\gamma+\frac{1}{2}e^{-2\gamma})}$ $=e^{n(\text{arccosh}(a)+a^{2}-a\sqrt{a^{2}-1}-\frac{1}{2})}.$
Proof of Proposition 3.14.
The approach is similar to the Laplace asymptotics, but this time we must be careful with lower-order terms, since we are looking for an upper bound that holds for all $n\in\mathbb{N}$. We begin (48) and deform the contour of integration from $\mathbb{R}$ to $w_{-}^{\ast}+\mathbb{R}$. Here, $w_{-}^{\ast}=i(a-\sqrt{a^{2}-1})$ is the critical point that lies on the contour $C_{-}$ we used for the Laplace asymptotics (in fact $C_{-}$ lies between $\mathbb{R}$ and $w_{-}^{\ast}+\mathbb{R}$); see Figure 7 for an illustration.
Figure 7: The contour used to prove the exponential bound for large $x$ values in Proposition 3.14
For convenience we write $u_{\ast}=a-\sqrt{a^{2}-1}$ so that $w_{-}^{\ast}=iu_{\ast}$. We obtain
$H\!\!\operatorname{e}_{n}(2a\sqrt{n})$ $=n^{n/2}\sqrt{\frac{n}{2\pi}}\,\int_{-\infty}^{\infty}\exp\left\{n\left[\log\left(2a+i\left(w+iu_{\ast}\right)\right)-\frac{(w+iu_{\ast})^{2}}{2}\right]\right\}\,\mathrm{d}w$
$=n^{n/2}\sqrt{\frac{n}{2\pi}}\,\intop_{-\infty}^{\infty}\exp\left\{n\left[\log\left(\frac{1}{u_{\ast}}+iw\right)-\frac{(w+iu_{\ast})^{2}}{2}\right]\right\}\,\mathrm{d}w$
where we have used the fact that $2a-u_{\ast}=a+\sqrt{a^{2}-1}=\frac{1}{u_{\ast}}$. Taking the absolute value inside the integral we obtain the inequality
$\left|H\!\!\operatorname{e}_{n}(2a\sqrt{n})\right|$ $\leq\frac{n^{(n+1)/2}}{\sqrt{2\pi}}\intop_{-\infty}^{\infty}\exp\left\{n\left(\frac{1}{2}\log\left(\frac{1}{u_{\ast}^{2}}+w^{2}\right)-\frac{w^{2}-u_{\ast}^{2}}{2}\right)\right\}\,dw$
Writing $\Phi(w)=\frac{1}{2}\log\left(\frac{1}{u_{\ast}^{2}}+w^{2}\right)-\frac{w^{2}-u_{\ast}^{2}}{2}$ and using the inequality $\log(1+x)\leq x$ for $x>0$:
$\Phi(w)=-\log u_{\ast}+\frac{1}{2}\log\left(1+w^{2}u_{\ast}^{2}\right)+\frac{u_{\ast}^{2}}{2}-\frac{w^{2}}{2}\leq\lambda_{1}(a)-\lambda_{2}(a)\frac{w^{2}}{2},$ (62)
where $\lambda_{1}(a)=-\log u_{\ast}+\frac{u^{2}_{\ast}}{2}$ and $\lambda_{2}(a)=1-u_{\ast}^{2}$. Crucially, $\lambda_{2}(a)>0$ since $u_{\ast}\in(0,1)$ for all $a>1$. More explicitly, $\lambda_{1}(a)=a^{2}-a\sqrt{a^{2}-1}-\frac{1}{2}+\text{arccosh}(a)$ and $\lambda_{2}(a)=2(1-a^{2}+a\sqrt{a^{2}-1})$. Putting everything together,
$\left|H\!\!\operatorname{e}_{n}(2a\sqrt{n})\right|$ $\leq\frac{n^{(n+1)/2}}{\sqrt{2\pi}}\,e^{n\lambda_{1}(a)}\intop_{-\infty}^{\infty}e^{-n\lambda_{2}(a)\frac{w^{2}}{2}}\,dw=\frac{n^{n/2}}{\sqrt{\lambda_{2}(a)}}\,e^{n\lambda_{1}(a)}$
as claimed. ∎
4 Applications to the oscillator wave functions $\Psi_{n}$ and the Hermite kernel $K_{n}$
@@ 4.1 Definitions of $\Psi_{n}$ and $K_{n}$
In Section 3.5 we saw that the Hermite polynomials go to zero exponentially fast outside the interval $2\sqrt{n},2\sqrt{n}]$. From this, we can deduce the behaviour of the wave oscillator functions $\Psi_{n}$.
@@
For all $a>1$ and all $n\in\mathbb{N}$,
$\Psi_{n}(2\sqrt{n}a)^{2}$ $\leq\frac{1}{4\pi\sqrt{n}(1-a^{2}+a\sqrt{a^{2}-1})}\exp\left\{2n\left[\text{arccosh}(a)-a\sqrt{a^{2}-1}\right]\right\}$
@@
Recall from Proposition 3.14 that
$\left|H\!\!\operatorname{e}_{n}(2\sqrt{n}a)\right|\leq\frac{n^{n/2}}{\sqrt{2(1-a^{2}+a\sqrt{a^{2}-1}})}\,e^{n(a^{2}-a\sqrt{a^{2}-1}-\frac{1}{2}+\text{arccosh}(a))}$
So by the inequality version of Stirling’s inequality, Remark 3.5,
$\Psi_{n}(2\sqrt{n}a)^{2}$ $\leq\frac{1}{\sqrt{2\pi}n!}e^{-2na^{2}}\frac{n^{n}}{2(1-a^{2}+a\sqrt{a^{2}-1})}e^{2n(a^{2}-a\sqrt{a^{2}-1}-\frac{1}{2}+\text{arccosh}(a))}$
$\leq\frac{1}{2\pi\sqrt{n}}\frac{1}{2(1-a^{2}+a\sqrt{a^{2}-1})}e^{2n\left[\text{arccosh}(a)-a\sqrt{a^{2}-1}\right]}$
as claimed. ∎
Remark 4.4.
Later on, specifically in the proof of Proposition 5.20, it will be useful to bound the rate of exponential decay in terms of a simpler function. To do so, note that the function $g$ defined by $g(a)=a\sqrt{a^{2}-1}-\text{arccosh}(a)$ satisfies $g(1)=0$ and also $g^{\prime}(a)=2\sqrt{a^{2}-1}>2(a-1)$ for $a>1$. Therefore $g(a)>(a-1)^{2}$ for all $a>1$. Moreover, the function $\lambda_{2}(a)=2(1-a^{2}+a\sqrt{a^{2}-1})$ is strictly increasing on $1,\infty)$ and satisfies $\lambda_{2}(1)=0$ . Therefore we can get the following bound: fix $\epsilon>0$ and let
$C_{\epsilon}$ $=\frac{1}{4\pi(2\epsilon-\epsilon^{2}+(1+\epsilon)\sqrt{2\epsilon-\epsilon^{2}}}.$ (65)
Then we have, for all $n\in\mathbb{N}$ and all $x>1+\epsilon$,
$\Psi_{n}\left(2\sqrt{n}a\right)^{2}$ $\leq\frac{C_{\epsilon}}{\sqrt{n}}e^{2n(\text{arccosh}(a)-a\sqrt{a^{2}-1})}\leq\frac{C_{\epsilon}}{\sqrt{n}}e^{-n(a-1)^{2}}.$ (66)
@@ 4.3 Cramèr’s bound for the global maximum
The function $\Psi_{n}(x)$ goes to zero exponentially fast outside of $x=[-2\sqrt{n},2\sqrt{n}]$, and the asymptotics of Section 3.4 suggest it should achieve its global maximum $\max_{x\in\mathbb{R}}|\Psi_{n}(x)|$ somewhere near the edge of this region $x=\pm 2\sqrt{n}$. However, since all of the asymptotics are proven pointwise, a different approach is needed in order to draw precise conclusions about the actual maximum. In [Indritz1961], a clever elementary analysis of the zeros of $\Psi^{\prime}_{n}$ reveals indeed that the maximum occurs at the closest zero of $\Psi_{n}^{\prime}$ to the edge $2\sqrt{n}$ and finds a bound for the maximum $\max_{x\in\mathbb{R}}|\Psi_{n}(x)|$. This bound is sometimes called Cramèr’s bound. We state this result here for later use.
@@
The proof is given in [Indritz1961]. Note that the function called $E_{n}(x)$ there is related to $\Psi_{n}(x)$ by $\Psi_{n}(x)=\frac{1}{2^{1/4}}E_{n}(x/\sqrt{2})$ . ∎
Remark 4.6.
In fact, the bound can be improved to yield
$cn^{-1/12}\leq\max_{x\in\mathbb{R}}\left|\Psi_{n}(x)\right|\leq Cn^{-1/12}$ (68)
for explicit constants $c,C>0$, see [Krasikov]. Both [Indritz1961] and [Krasikov] use the fact that for a given $n\in\mathbb{N}$, the supremum of $\Psi_{n}$ is after the last zero of $H\!\!\operatorname{e}_{n}$, near $2\sqrt{n}$, as can be seen in Figure 8.
@@ 4.4 Arcsine law limit for $\Psi_{n}^{2}$ and the limit of the quantum harmonic oscillator
By Proposition 4.3, $\Psi_{n}(\sqrt{n}x)\to 0$ for $|x|>2$ exponentially fast. To see the convergence to the function $\frac{1}{\pi\sqrt{4-x^{2}}}$ for $x\in[-2,2]$, we use the approximation (45) with $a=x/2$ to get
@@
The statement of the Schroedinger equation is exactly as in [griffithsQM], where the quantization for the energy is derived and it is shown that the solution is $\psi_{n}(x)=\left(\frac{m\omega}{\pi\hbar}\right)^{1/4}\frac{1}{\sqrt{2^{n}n!}}H_{n}(\xi)e^{-\xi^{2}/2}$ where $H_{n}$ are the physicists Hermite polynomials ($H_{1}(x)=2x,H_{2}(x)=4x^{2}-2$, etc) and $\xi=\sqrt{m\omega/\hbar}x$ is rescaled space. This is identical to our stated version by using the translation between physicists and probabilist $H_{n}(x)=2^{n/2}H\!\!\operatorname{e}_{n}(\sqrt{2}x)$ and noting that $cx=\sqrt{2}\xi$ by its definition. (One can also directly verify using the change of variable Section 7 that this solves the stated Schroedinger PDE)
@@ 4.5 Semi-circle law limit for $K_{n}$
The first statement follows by the simple change of variables $y=cx$ to establish that for any distribution $\rho$ and any constant $c$, one has:
@@
By the convergence of $\sqrt{n}\Psi_{n}(\sqrt{n}x)^{2}=\frac{\sqrt{n}}{\sqrt{2\pi}n!}H\!\!\operatorname{e}_{n}(\sqrt{n}x)^{2}e^{-\frac{1}{2}nx^{2}}$ to the arcsine law of Proposition 4.7, the compact support of the arcsine distribution on [2,2][-2,2] and the exponential decay of Proposition 4.3, we conclude that the moments of $\sqrt{n}\Psi_{n}(\sqrt{n}x)^{2}$ converge to the moments of the arcsine law.
Now, because the semi-circle law is the corresponding integral of the arcsine law from 70, by application of Lemma 4.11 we consequently see that the moments of $\sqrt{n}^{-1}K_{n}(\sqrt{n}x,\sqrt{n}x)$ converge to the moments of the semi-circle law. Since the semi-circle law is compactly supported on [2,2][-2,2], it is the unique distribution with these moments, and therefore the moment convergence implies the convergence in distribution (See e.g. Section 3.3.5 of [durrettPTE] on the moment problem and in particular Theorem 3.3.12) ∎
@@
The connection between the semi-circle law and the arcsine law has been noted in many other contexts too. In [HolcombVirag2019] the convergence is proven using random graphs. In [Ledoux2004], the convergence to the arcsine law is shown using differential equations. It is subsequently noted there that the mixture converges to $\sqrt{U}\xi_{arcsine}$ where $\xi_{arcsine}$ has the arcsine law, and $U$ is an independent uniform random variable. One can verify that $\sqrt{U}\xi_{arcsine}$ is distributed according to the semicircle law.